OpenStudy (anonymous):

Find a cubic function with the given zeros. 7, -3, 2

jimthompson5910 (jim_thompson5910):

If the given zeros are 7, -3, 2, then x = 7, -3, 2 x = 7, x = -3, x = 2 x-7=0, x+3 = 0, x-2=0 (x-7)(x+3)(x-2) = 0 I'll let you expand this and finish up

OpenStudy (anonymous):

Is it f(x) = x^3 - 6x^2 - 13x + 42?

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

nice job